# A Survey of Language-Based Communication in Robotics

### v1

@@ 5.2. Issues with LLMs in Robotic Systems
Some issues apply more specifically to the application of LLMs to robots. These are not unique to robotics, but are less likely to be a focus of the main LLM community, so should be considered separately to those in the previous section.
5.2.1. Modelling the Environment
@@
Modelling the world has been a crucial problem in robotics for decades(Brooks, 1991). Being essentially next-word prediction engines (“n-gram” models) (Brown et al., 1992), LLMs inherently do not include a representation of the environment and it can not easily be extended into their architecture. It has been argued that generalisation itself can only be effectively achieved through an understanding of the world which allows an agent to solve a new task in context (Mitchell, 2023; Mitchell et al., 2023). Understanding and planning at a human level will likely require the ability to reason hierarchically on multiple levels of abstraction and on different timescales (LeCun, 2022). While they can often prove effective in games and tasks where there is an imagined world, they tend to prove lackluster when asked to model an environment and track plans into the future. Examples such as stacking blocks are very difficult for LLMs to effectively solve (Valmeekam et al., 2024). Much work in this space aims to provide LLMs with a means of representing the environment that can be manipulated through text, the theory being that this would allow them to effectively understand the world and their effects upon it; a requirement that most deem important for an embodied robot when planning into the future (Jordanous, 2020). At the very least, it is necessary for a coherent language based plan which can be understood and approved by humans.
@@
A first-principles approach can be employed to this end which tests hypothesis about the environment directly in a physics engine (Liu et al., 2022b). This is likely too fundamental to be practical; more promising approaches maintain a textual description of the world state which is predicted, and then updated by the robot’s actual actions (Yoneda et al., 2023). Alternatively a simulation can be used as a world model (Xiang et al., 2024). The most common approach is to define a PDDL which the LLM can use to describe and update the environment (Silver et al., 2022, 2023). The key intuition here is that formal logic is text-based but provides a concrete and reliable way to model and understand the state of the world. Such an approach can effectively model relationships between objects and agents (Liu et al., 2023c; Guan et al., 2024). This can also be achieved with templated queries asking the LLM to describe within clearly defined expectations some relationship between objects. For example, “X is next to Y, how far should A be from B?” (Ding et al., 2023).
@@
An environment can also be imagined as an image; image generation models such as DALL-E can be used to create an image of what the goal state should be, and this is used to guide the model (Kapelyukh et al., 2023). Otherwise, LLMs can be used to resolve ambiguities in how objects might feasibly exist and be placed within an environment (Zhang et al., 2024b). This style of interpreting the environment visually may require more advanced models which can comprehend the nature of the physical world; in the same way transformers are suited to reasoning with language, similar architectures may prove useful in the domain of video (Selva et al., 2023), along with rich embeddings of video data (Sun et al., 2019). With recent developments in generative video models, it is possible that such architectures are on the horizon (Brooks et al., 2024). More complex relationships and environment understanding can also be modelling using representations like trees and graphs (Rana et al., 2023). Formalising a model of the environment in this way is sensible as such models are clear and reliable, the challenge is how the LLM queries and updates it; for such solutions to be practical there is need for strong handling of the interface between them (Chen et al., 2023c). One approach to this is to build two separate models, one classical LLM and another grounded model which interprets the environment. Their decisions can be combined in the same approximate means as SayCan (Brohan et al., 2023b), to perform the action which best satisfies both the overall plan, and the environmental state (Huang et al., 2024b).
@@
5.2.2. Memory and Nonstationarity
In dealing with both non-stationary, changing tasks and user demands, and potential damage and wear on the robot itself, some degree of adaption and lifelong learning may be required (Liu et al., 2021). This may consist of the agent adapting its policy, learning from its mistakes to become better at complex tasks, or even tuning its understanding of how to operate alongside humans (Irfan et al., 2021). Doing this would require a robot to update itself in one of two ways: (1) By finetuning its model constantly throughout its life, or (2) By storing a history or set of beliefs which are updated through experience. Finetuning may have issues around overfitting to the specific everyday tasks the robot is required to do (for example it may forget safety protocols which are not often needed). It also requires storing the complete LLM, unless a technique such as Low-Rank Adaptation is used (Hu et al., 2021). The option of storing history may be more promising, with LLMs emulating memory (Park et al., 2023; Bärmann et al., 2023). A wide variety of techniques are actively being explored, and are well described by Wang et al.. At a high-level though, three key issues arise: (1) What information to store, (2) How to store and summarise it, and (3) How to access and use those memories. Storing relevant information in the short term is already achieved with the standard context-window, although eventually memories will pass out of this “short-term memory” and will require a “long-term memory” system alongside this (Ng et al., 2023; Yao et al., 2023; Zhang et al., 2024a), and it can still be difficult for the LLM to pick out what is useful in this way (Kuratov et al., 2024). An extension of this idea is to summarise interactions by extracting the key sentiment and storing these as memories (Park et al., 2023), or to use embeddings or vector representations to capture some essence of the memory (Zhong et al., 2024). Such memories can be of an event, entity, or even a skill (Wang et al., 2023c). Once a database of memories has been collated, access can be achieved through Retrieval-Augmented Generation instead of feeding it into the main context window (Gao et al., 2023). Some degree of compression and forgetting may be required to maintain a useful record which is not bloated with information, much as humans do (Hou et al., 2024), or else careful consideration should be given of which information the LLM has already internalised to minimise unnecessary access or saving of information (Huang et al., 2024a).
@@ 5.2.2. Memory and Nonstationarity
5.2.3. Uncertainty of LLMs and Integrating them with other Models
@@
LLMs are inherently probabilistic, and so introduce uncertainty into their outputs (Adam and Carter, 2023), this is known as aleatoric uncertainty — the output of the model varies each time it is run due to a a variation in the selected next token. This is not uncommon in machine learning models, however the sequential nature of language (or a language-like series of actions) means that the small uncertainties compound and result in outputs increasingly divergent from what may be expected from the initial inputs (Ling et al., 2024). In a more practical sense, slight variations or unpredictabilities in a plan at the nascent stage are then built upon and a robot’s actions at a later point might be drastically different from what the designers intended. It is also important to consider epistemic uncertainty — There can also be situations in which the knowledge of LLMs or their conception of the environment is incorrect (Felicioni et al., 2024). In this case it is more feasible to argue that such deficiencies can be corrected through other models or checking with humans. Some approaches use the LLM itself to evaluate its uncertainty, although this can be argued as moving the problem to another unreliable model (Tanneru et al., 2024). Other approaches incorporate more reliable other models alongside the LLM to give an estimate of confidence, such as conformal prediction (Wang et al., [n. d.], 2024b). This could potentially be used to reduce the level of uncertainty in the system and improve the predictability of a language-based robot.
@@
Regardless of the means of reducing the uncertainty, this presents a significant problem for LLMs in robotics. It is likely that the most effective way of harnessing the opportunities of LLMs is to use them for planning at a high-level, and potentially for handling unexpected inputs, after which the LLM is used in conjunction with other models, or as a component of a complex architecture. Relying on RL and other models for the end-level actions is likely to be the most practical (Shi et al., 2023; Hu et al., 2023), or else using LLMs to empower the training of RL (Chen et al., 2024a) or create a step-by-step plan to be solved with RL (Shukla et al., 2023; Dalal et al., 2024). Language can act as the high-level operator of a set of APIs and other models which can be called upon like “tools” (Shen et al., 2024a; Qin et al., 2023; Schick et al., 2024).
5.2.4. Computation Speed and Memory Size
@@
LLMs are inherently slow to generate responses. Although parallelism is often employed during training it is not practical to distribute next-word prediction across different devices (Brakel et al., 2024). Many LLM-based robotics approaches operate at reasonably low frequencies (1-10Hz) (Brohan et al., 2023a), whereas practical robots are likely to require a frequency much higher (in the region of 500Hz) (Chignoli et al., 2021); it is not practical to run the LLM at anything close to this frequency for computational power reasons. Further to this, LLMs generally require a lot of memory to store, and are often not available to the public except through an API, which incurs additional latency. It is unlikely that language could practically be used for split-second decision-making, where simpler reactive policies are likely to be more reliable. Further to this, an API-based approach requires a constant online connection, something which is again impractical in many use cases, especially those where robots may be operating in hazardous environments with poor infrastructure.
5.2.5. Lack of Robot Datasets
@@
Datasets are often a limitation for robotics as gathering data can be expensive and time-consuming. LLMs may solve this issue, both through the virtue of training on a broad set of data which is generally freely available online (Liu et al., 2024a), and by generating their own training data using the LLM itself. This can take the form of self generated instructions to broaden the range of training data based on a few human-generated inputs (Wang et al., 2022). Alternatively, human responses can be used to identify bad generations and used to expand the dataset on which to train (Hancock et al., 2019), or to reinforce the intended role or expected style of the agent (Bae et al., 2022). Human labelling can also be used to identify if an image represents a success in a textually-defined task, which may be useful in training models to do this (Nair et al., 2022). Despite this, there are currently limitations on the amount of data and the number of simulation options available to developers of language-based robotics in certain cases (Liu et al., 2024b), and more mature options would be useful in future.
5.3. Ethical and Safety Concerns
@@ 5.3.1. Safety
AI safety as a concept predates most modern AI technologies, having been discussed at length in science fiction at least as early as Asimov’s famous three laws of robotics (Asimov, 1941). While these laws neither function particularly well in his writings, nor are practical to implement in most cases, their general theme persists in much of the AI safety literature. It should be noted though that much of this research is not responding to an Armageddon-like threat (McCauley, 2007), rather subtle issues such as bias and fairness, or lack of alignment with human morals when lives are at stake (Leslie, 2019). Some have employed inter-agent debate with human judgement as a strategy to deduce human moral codes (Irving et al., 2018). Generally, however, many concrete problems exist which are difficult to overcome (Amodei et al., 2016). Language models need only emulate human morals in the domain of text, which in some senses may be easier than doing so in action.
@@
The problem aligning LLMs to human morals is an active area of research (Wang et al., 2023e). Firstly, defining human morals is incredibly difficult. Some have resorted to language-based constitutions, which are derived from Asimov’s laws themselves (Ahn et al., 2024). Others have attempted to codify morals with Linear Temporal Logic (Yang et al., 2023b). This can either be injected into the system prompt before generation of text (Hua et al., 2024) or afterwards as a form of post-plan inspection (Yuan et al., 2024a) which can then be used to reject or alter the output (Liu et al., 2022a). As morals are so difficult to define, many have turned to humans to provide them. using a crowd-sourced labelling effort, models can be trained to inherently understand morals without a set of rules to follow (Ji et al., 2024), although this is likely to be less trustworthy in practise. One other aspect of safe and trustworthy language models is the issue of multilingual interface; LLMs generally operate best on the English language as that is what predominates their training data. Many are trained to have some ability in other languages but this tends to be limited (Kassner et al., 2021; Guo et al., 2020), and safety measures are often less effective on “low-resource” languages that the model has less training on (Shen et al., 2024b). It would be prudent to ensure those who interact in other languages have a similar experience and ability in order to ensure fair and trustworthy use by a wide variety of people (Chen et al., 2023b).
@@
Some efforts have sought to use datasets of safe and unsafe inputs and outputs to modify existing models such that they align better with human ethics. One method is to use the LLM itself to generate this data and further train itself (Xu et al., 2023a; Wang et al., 2022), although the scope of this is unlikely to be reliable in all cases. Alternatively humans can be inserted into a training loop for an LLM (Ouyang et al., 2022). In real-world use cases, guarantees of safety are unlikely to be strong, and when language-base robots inevitably make mistakes the most important outcome is that they can be learned from. Given an existing model, just a few safety examples can be used to finetune it (Yang et al., 2023c), although a careful balance must be pursued to ensure the model is not “paralysed” by safety (Bianchi et al., 2023). A model can also be pruned to reduce its size (Sun et al., 2023) which retains most of the power, and can increase safety (Hasan et al., 2024), although the consequential reduction in ability may make this impractical. Simple approaches to safety in language-based robotics limit the parameters of the output such as maximum acceleration (Yoshikawa et al., 2023), some define instinct-level behaviour to protect the robot and surrounding people (Zhang, 2023) while others incorporate a complex set of handwritten rules governing the entire output (Ahn et al., 2024). Regardless of the efficacy of safety approaches, trust will likely always be an issue with any autonomous system; people may be reluctant to rely upon a system they do not understand unless they are given appropriate evidence and experience that it is reliable (Zhou et al., 2024; Tekkesinoglu et al., 2024)
@@
For language-based robots to be trustworthy it is likely that a combination of handwritten rules and learning from mistakes will have to be employed. There is a substantial issue with public trust in either case, and research which provides reliable guarantees of safety would be highly impactful. Regardless, safety concerns are likely always to be relevant in the context of operating a large or unconstrained robot, especially in the proximity of humans.
5.3.2. Trust, Responsibility, and Governance

### v3

@@ 5.2 Issues with LLMs in Robotic Systems
Some issues apply more specifically to the application of LLMs to robotics. These are not unique to robotics, but are less likely to be a focus of the main LLM community, so should be considered separately from those in the previous section.
5.2.1 Modelling the Environment
@@
Modelling the world has been a crucial problem in robotics for decades Brooks, (1991). Being essentially next-word prediction engines (“n-gram” models) Brown et al., (1992), LLMs inherently do not include a representation of the environment and it can not easily be extended into their architecture. It has been argued that generalisation itself can only be effectively achieved through an understanding of the world which allows an agent to solve a new task in context Mitchell, (2023); Mitchell et al., (2023). Understanding and planning at a human level will likely require the ability to reason hierarchically on multiple levels of abstraction and on different timescales LeCun, (2022). While they can often prove effective in games and tasks where there is an imagined world, they tend to prove lackluster when asked to model an environment and track plans into the future. Examples such as stacking blocks are very difficult for LLMs to effectively solve Valmeekam et al., (2024). Much work in this space aims to provide LLMs with a means of representing the environment that can be manipulated through text, the theory being that this would allow them to effectively understand the world and their effects upon it; a requirement that most deem important for an embodied robot when planning into the future Jordanous, (2020). At the very least, it is necessary for a coherent language-based plan which can be understood and approved by humans.
@@
A first-principles approach can be employed to this end which tests hypotheses about the environment directly in a physics engine Liu et al., 2022b . This is likely too fundamental to be practical. More promising approaches maintain a textual description of the world state which is predicted and then updated by the robot’s actual actions Yoneda et al., (2023). Alternatively, a simulation can be used as a world model Xiang et al., (2024). The most common approach is to define a PDDL which the LLM can use to describe and update the environment Silver et al., (2022, 2023). The key intuition here is that formal logic is text-based but provides a concrete and reliable way to model and understand the state of the world. Such an approach can effectively model relationships between objects and agents Liu et al., 2023a ; Guan et al., (2024). This can also be achieved with templated queries asking the LLM to describe within clearly defined expectations some relationship between objects. For example, “X is next to Y, how far should A be from B?” Ding et al., (2023).
@@
An environment can also be imagined as an image; image generation models such as DALL-E can be used to create an image of what the goal state should be, and this is used to guide the model Kapelyukh et al., (2023). Otherwise, LLMs can be used to resolve ambiguities in how objects might feasibly exist and be placed within an environment Zhang et al., 2024a . This style of interpreting the environment visually may require more advanced models which can comprehend the nature of the physical world; in the same way, transformers are suited to reasoning with language, similar architectures may prove useful in the domain of video Selva et al., (2023), along with rich embeddings of video data Sun et al., (2019). With recent developments in generative video models, it is possible that such architectures are on the horizon Brooks et al., (2024). More complex relationships and environment understanding can also be modelled using representations like trees and graphs Rana et al., (2023). Formalising a model of the environment in this way is sensible as such models are clear and reliable, the challenge is how the LLM queries and updates it; for such solutions to be practical there is a need for strong handling of the interface between them Chen et al., 2023a . One approach to this is to build two separate models, one classical LLM and another grounded model which interprets the environment. Their decisions can be combined in the same approximate means as SayCan Brohan et al., 2023b , to perform the action which best satisfies both the overall plan, and the environmental state Huang et al., 2024b .
@@
5.2.2 Memory and Nonstationarity
In dealing with both non-stationary, changing tasks and user demands, and the potential damage and wear on the robot itself, some degree of adaptation and lifelong learning may be required Liu et al., (2021). This may consist of the agent adapting its policy, learning from its mistakes to become better at complex tasks, or even tuning its understanding of how to operate alongside humans Irfan et al., (2021). Doing this would require a robot to update itself in one of two ways: (1) by finetuning its model constantly throughout its life, or (2) by storing a history or set of beliefs which are updated through experience. Finetuning may have issues around overfitting to the specific everyday tasks the robot is required to do (for example it may forget safety protocols which are not often needed). It also requires storing the complete LLM, unless a technique such as Low-Rank Adaptation (LORA) is used Hu et al., (2021). The option of storing history may be more promising, with LLMs emulating memory Park et al., (2023); Bärmann et al., (2023). A wide variety of techniques are actively being explored, and are well described by Wang et al., 2024b . At a high level though, three key issues arise: (1) what information to store, (2) how to store and summarise it, and (3) how to access and use those memories. Storing relevant information in the short term is already achieved with the standard context window, although eventually, memories will pass out of this “short-term memory” and will require a “long-term memory” system alongside this Ng et al., (2023); Yao et al., (2023); Zhang et al., 2024b . It can still be difficult for the LLM to pick out what is useful in this way Kuratov et al., (2024). An extension of this idea is to summarise interactions by extracting the key sentiment and storing these as memories Park et al., (2023), or to use embeddings or vector representations to capture some essence of the memory Zhong et al., (2024). Such memories can be of an event, entity, or even a skill Wang et al., 2023a . Once a database of memories has been collated, access can be achieved through Retrieval-Augmented Generation (RAG) instead of feeding it into the main context window Gao et al., (2023). Some degree of compression and forgetting may be required to maintain a useful record which is not bloated with information, much as humans do Hou et al., (2024), or else careful consideration should be given of which information the LLM has already internalised to minimise unnecessary access or saving of information Huang et al., 2024a .
@@ 5.2.2 Memory and Nonstationarity
5.2.3 Uncertainty of LLMs and Integrating them with other Models
@@
LLMs are inherently probabilistic, and so introduce uncertainty into their outputs Adam and Carter, (2023), this is known as aleatoric uncertainty — the output of the model varies each time it is run due to a variation in the selected next token. This is not uncommon in machine learning models, however, the sequential nature of language (or a language-like series of actions) means that the small uncertainties compound and result in outputs increasingly divergent from what may be expected from the initial inputs Ling et al., (2024). In a more practical sense, slight variations or unpredictabilities in a plan at the nascent stage are then built upon and a robot’s actions at a later point might be drastically different from what the designers intended. It is also important to consider epistemic uncertainty — There can also be situations in which the knowledge of LLMs or their conception of the environment is incorrect Felicioni et al., (2024). In this case, it is more feasible to argue that such deficiencies can be corrected through other models or checking with humans. Some approaches use the LLM itself to evaluate its uncertainty, although this can be argued as moving the problem to another unreliable model Tanneru et al., (2024). Other approaches incorporate more reliable models alongside the LLM to give an estimate of confidence, such as conformal prediction Wang et al., 2023b ; Wang et al., 2024a . This could potentially be used to reduce the level of uncertainty in the system and improve the predictability of a language-based robot.
@@
Regardless of the means of reducing the uncertainty, this presents a significant problem for LLMs in robotics. It is likely that the most effective way of harnessing the opportunities of LLMs is to use them for planning at a high-level, and potentially for handling unexpected inputs, after which the LLM is used in conjunction with other models, or as a component of a complex architecture. Relying on RL and other models for the end-level actions is likely to be the most practical Shi et al., (2023); Hu et al., (2023), or else using LLMs to empower the training of RL Chen et al., 2024b or create a step-by-step plan to be solved with RL Shukla et al., (2023); Dalal et al., (2024). Language can act as the high-level operator of a set of APIs and other models which can be called upon like “tools” Shen et al., 2024b ; Qin et al., (2023); Schick et al., (2024).
5.2.4 Computation Speed and Memory Size
@@
LLMs are inherently slow to generate responses. Although parallelism is often employed during training it is not practical to distribute next-word prediction across different devices Brakel et al., (2024). Many LLM-based robotics approaches operate at reasonably low frequencies (1-10Hz) Brohan et al., 2023a , whereas practical robots are likely to require a frequency much higher (in the region of 500Hz) Chignoli et al., (2021); it is not practical to run the LLM at anything close to this frequency for computational power reasons. Further to this, LLMs generally require a lot of memory to store, and are often not available to the public except through an API, which incurs additional latency. It is unlikely that language could practically be used for split-second decision-making, where simpler reactive policies are likely to be more reliable. Further to this, an API-based approach requires a constant online connection — something which is also impractical in many use cases, especially those where robots may be operating in hazardous environments with poor or damaged infrastructure.
5.2.5 Lack of Robot Datasets
@@
Datasets are often a limitation for robotics as gathering data can be expensive and time-consuming. LLMs may solve this issue, both through the virtue of training on a broad set of data which is generally freely available online Liu et al., 2024b , and by generating their own training data using the LLM itself. This can take the form of self-generated instructions to broaden the range of training data based on a few human-generated inputs Wang et al., (2022). Alternatively, human responses can be used to identify bad generations and used to expand the dataset on which to train Hancock et al., (2019), or to reinforce the intended role or expected style of the agent Bae et al., (2022). Human labelling can also be used to identify if an image represents a success in a textually-defined task, which may be useful in training models to do this Nair et al., (2022). Despite this, there are currently limitations on the amount of data and the number of simulation options available to developers of language-based robotics in certain cases Liu et al., 2024a , and more mature options would be useful in future.
5.3 Ethical and Safety Concerns
@@ 5.3.1 Safety
AI safety as a concept predates most modern AI technologies, having been discussed at length in science fiction at least as early as Asimov’s famous three laws of robotics Asimov, (1941). While these laws neither function particularly well in his writings, nor are practical to implement in most cases, their general theme persists in much of the AI safety literature. It should be noted, however, that much of this research is not responding to an Armageddon-like threat McCauley, (2007), rather subtle issues such as bias and fairness, or lack of alignment with human morals when lives are at stake Leslie, (2019). Some have employed inter-agent debate with human judgement as a strategy to deduce human moral codes Irving et al., (2018). Generally, however, many concrete problems exist which are difficult to overcome Amodei et al., (2016). Language models need only emulate human morals in the domain of text, which in some senses may be easier than doing so in action.
@@
The problem of aligning LLMs to human morals is an active area of research Wang et al., 2023e . Firstly, defining human morals is incredibly difficult. Some have resorted to language-based constitutions, which are derived from Asimov’s laws themselves Ahn et al., (2024). Others have attempted to codify morals with Linear Temporal Logic Yang et al., 2023c . This can either be injected into the system prompt before generation of text Hua et al., (2024) or afterwards as a form of post-plan inspection Yuan et al., 2024b which can then be used to reject or alter the output Liu et al., 2022a . As morals are so difficult to define, many have turned to humans to provide them. Using a crowd-sourced labelling effort, models can be trained to inherently understand morals without a set of rules to follow Ji et al., (2024), although this is likely to be less trustworthy in practice. One other aspect of safe and trustworthy language models is the issue of multilingual interface; LLMs generally operate best in the English language as that is what predominates their training data. Many are trained to have some ability in other languages but this tends to be limited Kassner et al., (2021); Guo et al., (2020), and safety measures are often less effective on “low-resource” languages that the model has less training data for Shen et al., 2024a . It would be prudent to ensure that those who interact in other languages have a similar experience and ability in order to ensure fair and trustworthy use by a wide variety of people Chen et al., 2023e .
@@
Some efforts have sought to use datasets of safe and unsafe inputs and outputs to modify existing models such that they align better with human ethics. One method is to use the LLM itself to generate this data and further train itself Xu et al., 2023b ; Wang et al., (2022), although the scope of this is unlikely to be reliable in all cases. Alternatively, humans can be inserted into a training loop for an LLM Ouyang et al., (2022). In real-world use cases, guarantees of safety are unlikely to be strong, and when language-based robots inevitably make mistakes then the most important outcome is that they can be learned from. Given an existing model, just a few safety examples can be used to finetune it Yang et al., 2023b , although a careful balance must be pursued to ensure the model is not “paralysed” by safety Bianchi et al., (2023). A model can also be pruned to reduce its size Sun et al., (2023) which retains most of the power and can increase safety Hasan et al., (2024), although the consequential reduction in ability may make this impractical. Simple approaches to safety in language-based robotics limit the parameters of the output such as maximum acceleration Yoshikawa et al., (2023), some define instinct-level behaviour to protect the robot and surrounding people Zhang and Lu, (2023) while others incorporate a complex set of handwritten rules governing the entire output Ahn et al., (2024). Regardless of the efficacy of safety approaches, trust will likely always be an issue with any autonomous system; people may be reluctant to rely upon a system they do not understand unless they are given appropriate evidence and experience that it is reliable Zhou et al., (2024); Tekkesinoglu et al., (2024).
@@
For language-based robots to be trustworthy, a combination of handwritten rules and learning from mistakes will likely have to be employed. There is a substantial issue with public trust in either case, and research which provides reliable guarantees of safety would be highly impactful. Regardless, safety concerns are likely always to be relevant in the context of operating a large or unconstrained robot, especially in the proximity of humans.
5.3.2 Trust, Responsibility, and Governance